OpenStudy (anonymous):

Find x if f(x) = 2x + 7 and f(x) = -1.

OpenStudy (igreen):

If f(x) = -1 then that means you just input '-1' for 'x'. \(f(x) = 2x + 7\) \(f(-1) = 2(-1) + 7\) \(f(-1) = -2 + 7\) \(f(-1) = 5\)

OpenStudy (igreen):

You can give medals by clicking 'Best Response'. @livia2

OpenStudy (skullpatrol):

f(-1) means you just input '-1' for 'x'. @iGreen

OpenStudy (anonymous):

F(x) = 2x+7 and f(x) = -1 That means -1 = 2x+7 i.e 2x = -1-7 x = -4

OpenStudy (igreen):

?? @skullpatrol

OpenStudy (igreen):

Ohhh wait.. f(x) = 2x + 7 -1 = 2x + 7 -2x - 1 = 7 -2x = 8 x = -4